### v1

@@ 2.5 Signal Processing Pipeline
The first step, filtering the data, was common to both pipelines: a 0.5 Hz highpass filter was applied to remove low-frequency drifts and DC offset, and a 50 Hz notch filter was applied to suppress line noise from the Stentrode recordings. Both filters were implemented as zero-phase FIR filters using a Hamming window, with forward-backward filtering. The filter order was automatically determined according to MNE-Python’s default firwin design method [34]. The following subsections discuss the later steps in each pipeline separately.
@@ 2.5.1 Feature Visualisation
The temporal, spectral, and time-frequency features were visualised using event-related potentials (ERPs), point-biserial correlation coefficients ($r^{2}$), and spectrograms, respectively. A brief description of each technique is provided below.
@@
Evoked responses, or event-related potentials (ERPs), represent the average neural activity across multiple trials, time-locked to a specific event, such as the onset of a cue or saccade. Since trial averaging primarily preserves low-frequency components, a 15 Hz low-pass zero-phase FIR filter was applied to smooth the ERP plots for visualisation. Additionally, individual trials were plotted to observe ERP patterns at the single-trial level.
@@
$r(x_{1},x_{2})=\frac{\sqrt{N_{1}N_{2}}}{N_{1}+N_{2}}\cdot\frac{\mathrm{mean}(x_{1})-\mathrm{mean}(x_{2})}{\mathrm{std}(x_{1}\cup x_{2})}$ (1)
where $x_{1}$ and $x_{2}$ represent the power spectral density values in a given frequency bin for each condition and $N_{1}$ and $N_{2}$ denote the number of trials. The value of $r$ can be positive or negative. Since we are only interested in the magnitude of the difference, regardless of direction, it is common practice to report the squared values ($r^{2}$) to ensure that all values are positive. An $r^{2}$ value of 1 indicates that the feature is perfectly correlated with the class, while an $r^{2}$ value close to 0 suggests that the feature is not discriminative for the two classes.
@@ 2.5.2 Saccade Classification
To extract time-domain features, each trial was divided into non-overlapping 100 ms segments and the mean, standard deviation, variance, kurtosis, and root-mean-square (RMS) values were computed for each segment. For frequency-domain feature extraction, the mean and standard deviation of spectral power were calculated in the delta (0.5-4 Hz), theta (4-8 Hz), alpha (8-12 Hz), and beta (12-30 Hz) bands using the multitaper method.
For classification, different algorithms were investigated and the random forest classifier was found to perform at least as well as or better than other classifiers. The decoding algorithms were mainly implemented using the MNE-Python [34] and Scikit-learn [42] libraries in Python. The xDAWN denoising algorithm accepted the ‘n_components’ argument, which yielded the best decoding results when set to 2 components. The output of xDAWN was standardised to the range [0, 1] using min-max scaling, after which the scaled features were fed into a random forest classifier for classification. Grid search was performed to optimise the ‘n_estimators’ and ‘max_depth’ hyperparameters of the random forest classifier.
@@ 3.1 Time Domain Features: ERP Analysis
The next step was to visualise the ERPs in different directions for both tasks using the most discriminative channels. Figure 4 shows the evoked responses of different events (P300, initial saccade, back saccade), for direction-specific (left, right, up, down) and direction-agnostic (any direction) trials. The ‘any direction’ column represents ERPs of data pooled across all directions. Panel (a) represents the ERP of channel 10 in the sensor space, while the remaining panels (b)-(f) correspond to the source space data. In each ERP subplot, the amplitude is plotted against time, with a grey vertical line at $t=0$ representing event onset (arrow, initial saccade, back saccade). ‘N’ represents the number of trials across which the ERP average was computed. Furthermore, the black bold line represents the ERP of the fixation trials, which serve as a baseline for comparison, whereas the bold blue, purple, and green lines represent ERPs of P300, initial saccade, and back saccade trials, with thin traces showing individual trials.
@@ 3.1.3 Saccadic Potentials in Free-Viewing Task:
Although only ERPs of the most prominent channels are discussed here, Figure 9 in A shows ERPs of other good channels in the source space. For the Visually-Guided Task, the ERPs of IC4 (Figure 9(a)) demonstrated cue-related potentials occurring approximately 400-650 ms after arrow appearance. Furthermore, the ERP of IC2 for the initial saccade (Figure 9(b)) lacked a consistent ERP component, whereas for the back saccade (Figure 9(c)), the ERP of IC2 appeared as a positive peak that began just after the start of the saccade and lasted for about 250-300 ms after back saccade onset. For the Free-Viewing task, IC5 showed a positive ERP for both initial and back saccades, which began just before the onset of the saccade for the initial saccade (Figure 9(d)) and on saccade onset for the back saccade (Figure 9(e)) and lasted for approximately 200-300 ms after saccade onset. ERPs in different directions and single-trial plots showed somewhat similar patterns for free-viewing saccades.
@@ 3.2 Frequency Domain Features: Biserial Correlation Coefficient
For the Visually-Guided Task, Figures 5(a), 5(b), and 5(c) show the $r^{2}$ scores for all channels (y-axis) for the conditions(arrow onset, initial saccade and back saccade, respectively), with IC2 and IC4 showing high $r^{2}$ scores (red coloured) compared to other channels. For the P300 trials (Figure 5(a)), the discriminative features appeared in IC2 and IC4, primarily in the 8-24 Hz range, with IC2 showing consistent scores across all four directions. In contrast, the initial saccades (Figure 5(b)) exhibited discriminative features in the lower frequency range (below 8 Hz), with IC4 being more consistent across directions. The back saccades (Figure 5(c)) showed relatively small $r^{2}$ scores compared to the other two conditions. As the disriminative frequency bands were different when comparing P300 versus fixation, against saccade versus fixation, this suggests the presence of unique spectral features for P300 and visually-guided saccades.
For the Free-Viewing Task, Figures 5(d) and 5(e) present the $r^{2}$ scores for the initial and back saccades, respectively. Identical $r^{2}$ patterns across different channels were observed for all directions and both types of saccades. IC7 gave the highest $r^{2}$ scores, followed by IC10, IC4, and IC12. The most discriminative frequencies were found within the 6–25 Hz range, with peak discrimination occurring between 12–24 Hz. A general observation was that free-viewing saccades provided higher $r^{2}$ scores compared to visually-guided saccades, as reflected in the scale limits: $r^{2}$ = 0.4 for the Free-Viewing Task and $r^{2}$ = 0.1 for the Visually-Guided Task. This difference may be attributed to the experimental structure: unlike the Visually-Guided Task, where each saccade trial was preceded by a fixation trial, all fixation trials in the Free-Viewing Task occurred together, followed by all saccade trials. This temporal separation might have made the discrimination between fixation and saccade trials easier in the Free-Viewing Task.
@@ 3.3 Time-Frequency Domain Features: Spectrograms
For the Visually-Guided Saccade Task, Figures 6(a), 6(b), and 6(c) show the spectrograms for IC4, time-locked to the appearance of the visual cue, the initial saccade, and the back saccade, respectively. In Figure 6(a), the P300 potentials appeared as a synchronisation in low frequencies (red region below 10 Hz) accompanied by a desynchronisation in the beta band (blue region, 12-30 Hz), approximately 400 ms after arrow onset. These spectrotemporal patterns were consistently observed for saccades in all directions. The spectrograms for the initial saccade (Figure 6(b)) showed activity resembling the P300 potentials observed in Figure 6(a), occurring prior to the saccade onset ($t=0$). This suggests that the initial saccade activity was confounded by overlapping P300 responses. Moreover, Figure 6(c) did not reveal any distinct or consistent patterns for the back saccade.
For the Free-Viewing Task, Figures 6(d) and 6(e) show the spectrograms of IC7 for the initial and back saccades, respectively. Strong low-frequency synchronisation (below 15 Hz) was observed across most directions, indicated by red regions centred around the black vertical line ($t=0$) marking saccade onset. This pattern suggests the presence of consistent, time-locked spectral features associated with the initiation of free-viewing saccades.
@@ 3.4 Saccade Classification
Figure 7 presents the results of decoding within a [0.50.5-0.5- 0.50.50.50.50.5] s interval around saccade onset using sensor space data. Panels (a), (b), and (c) display decoding performance for initial saccades, back saccades, and combined saccades, respectively. Performance was measured using AUC-ROC scores, with the red horizontal dashed line indicating chance-level performance (AUC = 0.5). The bar colours represented different task types: blue for within-session (WS) decoding of the Visually-Guided Task, and orange, green, and purple for Session 1, Session 2, and pooled sessions of the Free-Viewing Task. Cross-session (CS) decoding results are shown in grey and black: grey bars indicate training on Session 2 and testing on Session 1, while black bars indicate the reverse. Each subplot includes decoding results for both saccade onset classification (fixation vs. saccade) and saccade direction classification (4-class and pairwise binary), grouped separately and labelled along the x-axis.
@@
For reference, Figure 10 in A shows the source space decoding results, which were consistent with the sensor space trends. Source space decoding generally achieved similar or higher AUC scores for both onset and direction classification. Notably, for cross-session saccade onset classification, AUC scores exceeded 0.85 for all conditions (initial, back and combined saccade), indicating strong generalisation.
@@
We were also interested in assessing the specific contribution of the pre-saccade interval to overall decoding performance. This distinction is important because decoding based on post-saccade data is more likely to reflect activity related to eye muscle movements, whereas oculomotor BCIs aim to decode the neural correlates of saccade planning. To investigate this, decoding was performed separately on pre-saccade, post-saccade, and combined pre+post-saccadic neural activity, with the results for decoding source space data presented in Table 1. Unlike previous analyses, a shorter epoch duration of 0.5 seconds was used instead of 1 second to prevent overlap between pre-saccade and post-saccade intervals of adjacent saccade events in the Free-Viewing Task. The results show that decoding performance using the pre-saccade interval is comparable to, though slightly lower than, that achieved with the post-saccade interval or the combined pre+post interval.
Table 1: Comparison of AUC-ROC scores for decoding three different 500 ms intervals relative to initial saccade onset using sensor space data. The ‘pre+post’, ‘pre’, and ‘post’ intervals correspond to the time windows $[-250,250]$ ms, $[-500,0]$ ms, and $[0,500]$ ms, respectively, all time-locked to saccade onset ($t=0$). ‘S1’, ‘S2’ and ‘S1+S2’ refer to data from session 1, session 2, and the pooled data, respectively, of the Free-Viewing Task. ‘Test: Free-Viewing S1’ indicates training on Session 2 and testing on Session 1, while ‘Test: Free-Viewing S2’ refers to training on Session 1 and testing on Session 2.
@@ 4.1 Comparison of Saccadic Potentials between Visually-Guided and Free-Viewing Saccade Tasks
While the evoked responses of free-viewing saccades time-locked to saccade onset showed identical saccade-related potentials for each saccade direction (left, right, up, and down), they had a different time course than the visually-guided saccades, which was particularly true for the initial saccade in the Visually-Guided Task. This can be explained by the presence of P300 potentials in the Visually-Guided Task due to the appearance of visual cues (arrows onsets) that were absent in the Free-Viewing Task. These evoked responses overlap with individual trials differently; therefore, ERPs appear differently in different directions. The diverse impact of P300 potentials on individual trials can be explained by investigating the distribution of wait times between the onset of the cue (arrow) and the actual onset of the saccade. As shown in Figure 8, for valid visually-guided trials ($N~{}=~{}160$), the user moved their eyes on average 626 ms after arrow appearance (min = 414 ms, max = 935 ms, median = 612 ms, std = 103 ms). This explains why average ERPs appear differently in different directions due to the impact of P300 potentials. I seems to us that saccade-related ERPs follow cue-related ERPs; since P300 occurs approximately 300 ms after visual stimulus onset and the distribution of wait times suggests that their effect should be reduced around the onset of the saccades. However, it is difficult to separate saccadic potentials from P300 for the Visually-Guided Task as both features overlap in the time domain at low frequencies.
@@ 4.2 Potential for an Endovascular P300 BCI
4.3 Saccade Planning or EOG Artefacts?
An interesting question is whether the proposed oculomotor BCI system decodes neural activity related to the planning and execution of saccades or simply detects electrooculography (EOG) artefacts present in the recorded signals. EOG signals [22] are commonly observed in scalp EEG recordings and occur due to physical eye movements (corneo-retinal potentials). These artefacts are an order of magnitude larger than the underlying EEG signal and are particularly dominant in the frontal channels.
One way to address this ambiguity is by comparing the waveforms of saccadic potentials against typical EOG artefacts observed in the EEG data. As shown in Figure 4(e-f), these saccade-related potentials also include a pre-saccade component, which precedes physical eye movement and cannot be attributed to EOG artefacts. Furthermore, EOG artefacts observed in EEG recordings typically show opposite polarity in contralateral directions and differ between vertical and horizontal saccades (for details, see [50, 23, 25, 51]). In contrast, saccadic potentials appear identical across all saccade directions, suggesting they are linked to the planning and execution of saccades, regardless of direction. Furthermore, the EOG artefacts in the EEG data are significantly larger than actual brain signals and are most prominent in channels closer to the eyes. Since the Stentrode is implanted near the supplementary motor area, it is further from the source of the EOG artefacts. As a result, no such artefacts were observed during the visual inspection of the original data, similar to the recordings from other invasive neural implants.
An alternative approach is to repeat the decoding analysis using only the pre-saccade interval, and our results in Table 1 showed that decoding performance using the pre-saccade interval is comparable to, though slightly lower than, performance with the combined pre+post interval or the post-saccade interval. This difference can be attributed to the activity of the movement neurons that occurs both before and after saccade execution [52] and thus using both pre-saccade and post-saccade activity results in higher performance.

### v3

@@ 2.5 Signal Processing Pipeline
The first step, filtering the data, was common to both pipelines: a 0.5 Hz highpass filter was applied to remove low-frequency drifts and DC offset, and a 50 Hz notch filter was applied to suppress line noise from the Stentrode recordings. Both filters were implemented as zero-phase FIR filters using a Hamming window with forward-backward filtering. The filter order was automatically determined according to MNE-Python’s default ‘firwin’ method [30]. The following subsections discuss the later steps in each pipeline separately.
@@ 2.5.1 Feature Visualisation
The temporal, spectral, and time-frequency features were visualised using event-related potentials, point-biserial correlation coefficients ($r^{2}$), and spectrograms, respectively. A brief description of each technique is provided below.
@@
Evoked responses, or event-related potentials (ERPs), represent the average neural activity across multiple trials, time-locked to a specific event, such as the onset of a cue or saccade. Since trial averaging primarily preserves low-frequency components, a 15 Hz low-pass zero-phase FIR filter was applied to smooth the ERP plots for visualisation. In addition, individual trials were plotted to observe ERP patterns at the single-trial level.
@@
$r(x_{1},x_{2})=\frac{\sqrt{N_{1}N_{2}}}{N_{1}+N_{2}}\frac{\mathrm{mean}(x_{1})-\mathrm{mean}(x_{2})}{\mathrm{std}(x_{1}\cup x_{2})},$ (1)
where $x_{1}$ and $x_{2}$ represent the power spectral density values in a given frequency bin for each condition and $N_{1}$ and $N_{2}$ denote the numbers of trials. The value of $r$ can be positive or negative. Since we are only interested in the magnitude of the difference, regardless of direction, it is common practice to report the squared values ($r^{2}$) to ensure that all values are positive. An $r^{2}$ value of 1 indicates that the feature is perfectly correlated with the class, while an $r^{2}$ value close to 0 suggests that the feature is not discriminative for the two classes.
@@ 2.5.2 Saccade Classification
To extract time-domain features, each trial was divided into non-overlapping 100 ms segments and the mean, standard deviation, variance, kurtosis, and root-mean-square (RMS) values were computed for each segment. For frequency domain feature extraction, the mean and standard deviation of spectral power were calculated in the delta (0.5-4 Hz), theta (4-8 Hz), alpha (8-12 Hz), and beta (12-30 Hz) bands using the multitaper method.
For classification, multiple algorithms were evaluated, with the Random Forest classifier performing slightly better than the others (see A for details). To account for class imbalances resulting from the removal of bad trials, the results are reported using AUC-ROC scores rather than accuracy. The decoding algorithms were mainly implemented using the MNE-Python [30] and Scikit-learn [42] libraries in Python. The xDAWN denoising algorithm accepted the ‘n_components’ argument, which yielded the best decoding results when set to 2 components. The output of xDAWN was standardised to the range [0, 1] using min-max scaling, after which the scaled features were fed into a random forest classifier for classification. Grid search was performed to optimise the ‘n_estimators’ and ‘max_depth’ hyperparameters of the random forest classifier.
@@ 3.1 Time Domain Features: ERP Analysis
The next step was to visualise the ERPs in different directions for both tasks using the most discriminative channels. Figure 5 shows the evoked responses of different events (P300, initial saccade, back saccade), for direction-specific (left, right, up, down) and direction-agnostic (any direction) trials. The ‘any direction’ column represents ERPs of data pooled across all directions. Panel (a) represents the ERP of channel 10 in the sensor space, while the remaining panels (b)-(f) correspond to the source space data. In each ERP subplot, the amplitude is plotted against time, with a grey vertical line at $t=0$ representing event onset (arrow, initial saccade, back saccade). ‘N’ represents the number of trials across which the ERP average was computed. Furthermore, the black bold line represents the ERP of the fixation trials, which serve as a baseline for comparison, whereas the bold blue, purple, and green lines represent ERPs of P300, initial saccade, and back saccade trials, with thin traces showing individual trials. It is worth noting that the ERPs in different panels are plotted on different timescales, chosen to ensure that the ERP components return to baseline while remaining short enough to allow visualisation of features in individual trials.
@@ 3.1.3 Saccadic Potentials in Free-Viewing Task:
Although only ERPs of the most prominent channels are discussed here, Figure 11 in B shows ERPs of other good channels in the source space. For the Visually-Guided Task, the ERPs of IC4 (Figure 11(a)) demonstrated cue-related potentials occurring approximately 400-650 ms after arrow appearance. Furthermore, the ERP of IC2 for the initial saccade (Figure 11(b)) lacked a consistent ERP component, whereas for the back saccade (Figure 11(c)), the ERP of IC2 appeared as a positive peak that began just after the start of the saccade and lasted for about 250-300 ms after back saccade onset. For the Free-Viewing task, IC5 showed a positive ERP for both initial and back saccades, which began just before the onset of the saccade for the initial saccade (Figure 11(d)) and on saccade onset for the back saccade (Figure 11(e)) and lasted for approximately 200-300 ms after saccade onset. ERPs in different directions and single-trial plots showed somewhat similar patterns for free-viewing saccades.
@@ 3.2 Frequency Domain Features: Biserial Correlation Coefficient
For the Visually-Guided Task, Figures 6(a), 6(b), and 6(c) show the $r^{2}$ scores for all channels (y-axis) for the conditions(arrow onset, initial saccade and back saccade, respectively), with IC2 and IC4 showing high $r^{2}$ scores (red coloured) compared to other channels. For the P300 trials (Figure 6(a)), the discriminative features appeared in IC2 and IC4, primarily in the 8-24 Hz range, with IC2 showing consistent scores across all four directions. In contrast, the initial saccades (Figure 6(b)) exhibited discriminative features in the lower frequency range (below 8 Hz), with IC4 being more consistent across directions. The back saccades (Figure 6(c)) showed relatively small $r^{2}$ scores compared to the other two conditions. As the disriminative frequency bands were different when comparing P300 versus fixation, against saccade versus fixation, this suggests the presence of unique spectral features for P300 and visually-guided saccades.
For the Free-Viewing Task, Figures 6(d) and 6(e) present the $r^{2}$ scores for the initial and back saccades, respectively. Identical $r^{2}$ patterns across different channels were observed for all directions and both types of saccades. IC7 gave the highest $r^{2}$ scores, followed by IC10, IC4, and IC12. The most discriminative frequencies were found within the 6–25 Hz range, with peak discrimination occurring between 12–24 Hz. A general observation was that free-viewing saccades provided higher $r^{2}$ scores compared to visually-guided saccades, as reflected in the scale limits: $r^{2}$ = 0.4 for the Free-Viewing Task and $r^{2}$ = 0.1 for the Visually-Guided Task. This difference may be attributed to the experimental structure: unlike the Visually-Guided Task, where each saccade trial was preceded by a fixation trial, all fixation trials in the Free-Viewing Task occurred together, followed by all saccade trials. This temporal separation might have made the discrimination between fixation and saccade trials easier in the Free-Viewing Task.
@@ 3.3 Time-Frequency Domain Features: Spectrograms
For the Visually-Guided Saccade Task, Figures 7(a), 7(b), and 7(c) show the spectrograms for IC4, time-locked to the appearance of the visual cue, the initial saccade, and the back saccade, respectively. In Figure 7(a), the P300 potentials appeared as a synchronisation in low frequencies (red region below 10 Hz) accompanied by a desynchronisation in the beta band (blue region, 12-30 Hz), approximately 400 ms after arrow onset. These spectrotemporal patterns were consistently observed for saccades in all directions. The spectrograms for the initial saccade (Figure 7(b)) showed activity resembling the P300 potentials observed in Figure 7(a), occurring prior to the saccade onset ($t=0$). This suggests that the initial saccade activity was confounded by overlapping P300 responses. Moreover, Figure 7(c) did not reveal any distinct or consistent patterns for the back saccade.
For the Free-Viewing Task, Figures 7(d) and 7(e) show the spectrograms of IC7 for the initial and back saccades, respectively. Strong low-frequency synchronisation (below 15 Hz) was observed across most directions, indicated by red regions centred around the black vertical line ($t=0$) marking saccade onset. This pattern suggests the presence of consistent, time-locked spectral features associated with the initiation of free-viewing saccades.
@@ 3.4 Saccade Classification
Figure 8 presents the results of decoding within a [0.5-0.50.50.5] s interval around saccade onset using sensor space data. Panels (a), (b), and (c) display decoding performance for initial saccades, back saccades, and combined saccades, respectively. Performance was measured using AUC-ROC scores, with the red horizontal dashed line indicating chance-level performance (AUC = 0.5). The bar colours represented different task types: blue for within-session (WS) decoding of the Visually-Guided Task, and orange, green, and purple for Session 1, Session 2, and pooled sessions of the Free-Viewing Task. Cross-session (CS) decoding results are shown in grey and black: grey bars indicate training on Session 2 and testing on Session 1, while black bars indicate the reverse. Each subplot includes decoding results for both saccade onset classification (fixation vs. saccade) and saccade direction classification (4-class and pairwise binary), grouped separately and labelled along the x-axis.
@@
For reference, Figure 12 in C shows the source space decoding results, which were consistent with the sensor space trends. Source space decoding generally achieved similar or higher AUC scores for both onset and direction classification. Notably, for cross-session saccade onset classification, AUC scores exceeded 0.85 for all conditions (initial, back and combined saccade), indicating strong generalisation.
@@
We also assessed the specific contribution of the pre-saccade interval to overall decoding performance. This distinction is important because decoding based on post-saccade data is more likely to reflect activity related to eye muscle movements, whereas oculomotor BCIs aim to decode the neural correlates of saccade planning. To investigate this, decoding was performed separately on pre-saccade, post-saccade, and combined pre+post-saccadic neural activity, with the results for decoding source space data presented in Table 1. Unlike previous analyses, a shorter epoch duration of 0.5 s was used instead of 1 s to prevent overlap between pre-saccade and post-saccade intervals of adjacent saccade events in the Free-Viewing Task. The results show that decoding performance using the pre-saccade interval is comparable to, though slightly lower than, that achieved with the post-saccade interval or the combined pre+post interval.
Table 1: Comparison of AUC-ROC scores for decoding three different 500 ms intervals relative to initial saccade onset using sensor space data. The ‘pre+post’, ‘pre’, and ‘post’ intervals correspond to the time windows [250,250][-250,250] ms, [500,0][-500,0] ms, and $[0,500]$ ms, respectively, all time-locked to saccade onset ($t=0$). ‘S1’, ‘S2’ and ‘S1+S2’ refer to data from session 1, session 2, and the pooled data, respectively, of the Free-Viewing Task. ‘Test: Free-Viewing S1’ indicates training on Session 2 and testing on Session 1, while ‘Test: Free-Viewing S2’ refers to training on Session 1 and testing on Session 2.
@@ 4.1 Comparison of Saccadic Potentials between Visually-Guided and Free-Viewing Saccade Tasks
While the evoked responses of free-viewing saccades time-locked to saccade onset showed identical saccade-related potentials for each saccade direction (left, right, up, and down), they had a different time course than the visually-guided saccades, which was particularly true for the initial saccade in the Visually-Guided Task. This can be explained by the presence of P300 potentials in the Visually-Guided Task due to the appearance of visual cues (arrows onsets) that were absent in the Free-Viewing Task. These evoked responses overlap with individual trials differently; therefore, ERPs appeared differently in different directions. The diverse impact of P300 potentials on individual trials can be explained by investigating the distribution of wait times between the onset of the cue (arrow) and the actual onset of the saccade. As shown in Figure 9, for valid visually-guided trials (N=160N~=~160), the user moved their eyes on average 626 ms after arrow appearance (min = 414 ms, max = 935 ms, median = 612 ms, std = 103 ms). This explains why average ERPs appeared differently in different directions due to the impact of P300 potentials. It appeared that saccade-related ERPs follow cue-related ERPs, since P300 occurs approximately 300 ms after visual stimulus onset and the distribution of wait times suggests that their effect should be reduced around the onset of the saccades. However, it is difficult to separate saccadic potentials from P300 for the Visually-Guided Task as both features overlap in the time domain at low frequencies.
@@ 4.2 Potential for an Endovascular P300 BCI
4.3 Neural Correlates or EOG Artefacts?
An important question is whether the proposed oculomotor BCI system decodes neural activity related to the planning of saccades, or whether it merely detects EOG artefacts that are commonly observed in scalp EEG recordings [50]. We present several reasons to support the interpretation that the saccade-related potentials observed in our study are neural in origin and cannot be attributed to EOG artefacts.
First, visual inspection of both the original sensor space data and the source space data (after applying ICA) did not reveal periodic waveforms typically associated with EOG artefacts in EEG. Second, the Stentrode is placed in the superior sagittal sinus, which is anatomically distant from the corneo-retinal dipole. This location is more analogous to subdural ECoG arrays, which are generally unaffected by EOG artefacts. Third, the saccade-related potentials shown in Figure 5(e–f) include a prominent pre-saccade component that precedes physical eye movement. Since EOG artefacts arise from actual eye motion, they cannot account for this pre-movement activity. Fourth, EOG artefacts in EEG typically show polarity reversal across contralateral directions and differ substantially between horizontal and vertical saccades [51, 52, 53, 54]. In contrast, the saccade-related potentials recorded in our study appear consistent across all directions, suggesting a common underlying neural process rather than direction-dependent EOG artefacts. Finally, the strongest evidence comes from our decoding analysis. As shown in Table 1, the classification performance using only pre-saccade data is comparable to performance using both pre-saccade and post-saccade windows. Since EOG artefacts cannot precede eye movement onset, this indicates that the classifier relies on pre-movement neural activity rather than ocular artefacts.